OpenStudy (anonymous):

can one of yall help me with thi s...y+2/7=3x?????

OpenStudy (anonymous):

so? what do want with it?

OpenStudy (anonymous):

what are you looking for? slope? y intercept? graph?

OpenStudy (anonymous):

she wants to find slope...posted it minutes back...

OpenStudy (anonymous):

\[\frac{y+2}{7}=3x?\]

OpenStudy (anonymous):

is this what u have? or y+ (2/7)=3x